# Mixed Near-field and Far-field Target Localization for Low-altitude Economy

### v1

@@ II-B Signal Model
The $K$ UAVs are assumed to simultaneously transmit pilots in $M$ time slots which are in narrow-band and uncorrelated. As such, the received signal in the UPA antennas at time $m$, denoted by $\bm{\eta}(m)\in\mathbb{C}^{N\times 1}$, is given by
@@
where $\mathbf{G}\triangleq[\mathbf{A},\mathbf{B}]\in\mathrm{C}^{N\times K}$ denotes the array steering matrix (ASM) with $\mathbf{A}\triangleq[\mathbf{a}\left(\theta_{1},\phi_{1}\right),\mathbf{a}\left(\theta_{2},\phi_{2}\right),\cdots,\mathbf{a}\left(\theta_{K_{1}},\phi_{K_{1}}\right)]$ and $\mathbf{B}\triangleq[\mathbf{b}\left(r_{K_{1}+1},\theta_{K_{1}+1},\phi_{K_{1}+1}\right),\cdots,\mathbf{b}\left(r_{K},\theta_{K},\phi_{K}\right)]\in\mathbb{C}^{N\times K_{2}}$ denoting the ASM for the far- and near-field UAVs, respectively. The noise vector is distributed as $\mathbf{n}(m)\sim\mathcal{C\mathcal{N}}\left(0,\sigma^{2}\mathbf{I}_{M}\right)$ with zero mean and variance $\sigma^{2}$. Moreover, $\mathbf{s}(m)=[\mathbf{s}^{T}_{\rm F}(m),\mathbf{s}^{T}_{\rm N}(m)]^{T}\in\mathbb{C}^{K\times 1}$ denotes the source vector (i.e., transmitted signals by UAVs) with $\mathbf{s}_{\rm F}(m)=[s_{1}(m),s_{2}(m),\cdots,s_{K_{1}}(m)]^{T}\in\mathbb{C}^{K_{1}\times 1}$ and $\mathbf{s}_{\rm N}(m)=[s_{K_{1}+1}(m),s_{K_{1}+2}(m),\cdots,s_{K}(m)]^{T}\in\mathbb{C}^{K_{2}\times 1}$ representing the source sub-vectors for far- and near-filed UAVs, satisfying
$\mathbb{E}\{s_{j}(m)s_{k}(m)\}=\left\{\begin{matrix}g_{k},\leavevmode\nobreak\ \leavevmode\nobreak\ j=k,\\ 0,\leavevmode\nobreak\ \leavevmode\nobreak\ j\neq k,\end{matrix}\right.$ (9)
where $p_{k}$ and $g_{k}=N|\beta_{k}|^{2}p_{k}$ denote the transmit power and equivalent channel power gain for the $k$-th UAV, respectively.
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ =\mathbf{W}(m)\mathbf{A}\mathbf{s}_{{\rm F}}(m)\!\!+\!\!\mathbf{W}(m)\mathbf{B}\mathbf{s}_{{\rm N}}(m)\!\!+\!\!\mathbf{W}(m)\mathbf{n}(m),$
@@
where $\bm{\Pi}\in\mathbb{C}^{N\times N}$ and $\mathbf{v}_{f}(m)\in\mathbb{C}^{U\times 1},f\in\mathcal{F}\triangleq\{1,2,\cdots,N_{\rm RF}\}$ denote the permutation matrix determined by the sub-connected configuration and the combining vector of the $f$-th RF chain, respectively. For example, consider a UPA with $N_{x}=2$ and $N_{y}=4$ with $N_{\rm RF}=2$ RF chains. Following the antennas’ indices, the combining vectors of the first and second RF chains are given by $[\mathbf{W}^{T}(m)]_{:,1}=[v_{1,1}(m),v_{1,2}(m),0,0,v_{1,3}(m),v_{1,4}(m),0,0]^{T}$, and $[\mathbf{W}^{T}(m)]_{:,2}=[0,0,v_{2,1}(m),v_{2,2}(m),0,0,v_{2,3}(m),v_{2,4}(m)]^{T}$, respectively. In addition, the permutation matrix $\bm{\Pi}$ can be constructed as $\bm{\Pi}=[\mathbf{e}_{1},\mathbf{e}_{2},\mathbf{e}_{5},\mathbf{e}_{6},\mathbf{e}_{3},\mathbf{e}_{4},\mathbf{e}_{7},\mathbf{e}_{8}]^{T}$, where $\mathbf{e}_{n}$ denotes the $n$-th column of the identity matrix $\mathbf{I}_{8}$. Hence, the combining matrix $\mathbf{W}^{T}(m)$ can be rewritten as $\mathbf{W}^{T}(m)=\bm{\Pi}\times\text{diag}\{\mathbf{v}_{1}(m),\mathbf{v}_{2}(m)\}$, where $\mathbf{v}_{1}(m)=[v_{1,1}(m),v_{1,2}(m),v_{1,3}(m),v_{1,4}(m)]^{T}$ and $\mathbf{v}_{2}(m)=[v_{2,1}(m),v_{2,2}(m),v_{2,3}(m),v_{2,4}(m)]^{T}$.
@@ III Proposed Mixed-field Localization Method
In this section, we first elaborate the main design challenges when the BS of a hybrid antenna array is used for localizing the mixed-field UAVs. Then an efficient localization algorithm is proposed to solve them.
Specifically, compared with existing works on mixed-field localization, the studied problem differs in the following three main aspects, which introduce new challenges. First, most existing works on mixed-field localization (e.g., [16]) rely on the assumption that the inter-antenna spacing is no greater than $\frac{\lambda}{4}$, such that reduced-dimensional/modified MUSIC algorithms can be devised to enable near-field localization. However, this assumption is incompatible to legacy wireless communication systems with typically half-wavelength inter-antenna spacing. Second, existing angle/location estimation methods primarily focus on digital array systems, which, however, may not be applicable to hybrid antenna systems with reduced-dimensional received signals. [^1] Although the beamspace MUSIC method was designed to effectively estimate path directions hybrid antenna systems [24], it inevitably decreases the maximum number of estimable sources $\mathcal{O}(N_{\rm RF})$. Third, existing mixed-field localization methods differentiate far-field and near-field targets by detecting far-field targets’ angles in the angle-domain MUSIC spectrum by letting $r\to\infty$. However, it will be shown later in Section III-C that this method may make false detection of near-field UAVs, when their ranges are relatively large (still smaller than Rayleigh distance), hence resulting in high localization errors.
@@
  • In Step 1, we accurately recover equivalent received signals at the UPA antennas from the reduced-dimensional signals $\mathbf{y}(m)$, by judiciously designing the analog combining matrices over time $\{\mathbf{W}(m)\}$. In particular, the UPA with half-wavelength inter-antenna spacing is considered, which inevitably cause angular ambiguity and will be addressed in the third step.
  • In Step 2, we design a modified low-dimensional MUSIC algorithm for UPA-based target localization by extracting the anti-diagonal entries of original received covariance matrix. Based on this, we further construct a new Toeplitz-structured (far-field equivalent) covariance matrix, which enables effective decoupling of the angle and range estimation.
  • In Step 3, we propose an efficient UAV classification method based on distinct characteristics of 3D MUSIC spectrum in the range domain between far-field and near-field UAVs, which can also effectively address the angular ambiguity issue caused by half-wavelength inter-antenna spacing.
@@ III-A Signal Reconstruction
To this end, we divide the $M$ received signals $\mathbf{y}(m)$ into $L=M/U$ groups, for which the $\ell$-th group’s signals can be concatenated into a new effective signal vector $\mathbf{Y}_{\ell}\in\mathbb{C}^{N\times 1}$ as given in (12). As such, each group’s received pilots can be utilized to recover a sample of equivalent signals received at the UPA. [^2] In (12), for ease of design, the transmitted signals from UAVs are assumed to be identical in every $M$ pilots, i.e., $\mathbf{s}\big{(}M(\ell-1)+1\big{)}=\mathbf{s}\big{(}M(\ell-1)+2\big{)}=,\cdots,=\mathbf{s}\big{(}M(\ell-1)+M\big{)}\triangleq\mathbf{s}_{\ell}$. Moreover, for simplicity, we set $\mathbf{W}(u)=\mathbf{W}\big{(}u+\ell U\big{)},\forall u\in\mathcal{U}\triangleq\{1,2,\cdots,U\},\forall\ell\in\mathcal{L}$. As such, if ${\mathbf{W}_{0}}\triangleq[\mathbf{W}^{T}(1),\mathbf{W}^{T}(2),\cdots,\mathbf{W}^{T}(U)]^{T}$ is an invertible matrix, the $\ell$-th sample of the reconstructed signals received at the UPA antennas can be recovered as
@@
$({\bf P1}):\leavevmode\nobreak\ \leavevmode\nobreak$ $\min_{\mathbf{W}_{0}}\leavevmode\nobreak\ {\rm Tr}\Big{(}\mathbf{n}_{\ell}\mathbf{n}_{\ell}^{H}\Big{)}$ (15a)
$\leavevmode\nobreak\ {\text{s}}{\text{.t}}{\rm{.}}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\rm Rank}(\mathbf{W}_{0})=N,$ (15b)
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ |[\mathbf{v}_{f}(u)]_{i}|=1,\forall u,i\in\mathcal{U},\forall f\in\mathcal{F},$ (15c)
@@
Although Problem (P1) is non-convex, we provide the necessary conditions for its optimal solution below by using similar solution methods in [25].
@@ III-B Angle Estimation
In this subsection, we first decouple the joint angle-and-range estimation into separate angle and range estimation. Then, we construct a modified covariance matrix (determined by targets’ angles only) and devise a decoupled 2D-angular MUSIC algorithm to estimate candidate angles for both near-field and far-field UAVs.
@@
where $\mathbf{S}_{\rm F}=\mathbb{E}\left\{\mathbf{s}_{\rm F}(t)\mathbf{s}_{\rm F}^{\mathrm{H}}(t)\right\}=\text{diag}\{g_{1},g_{2},\cdots,g_{K_{1}}\}$ and $\mathbf{S}_{\rm N}=\mathbb{E}\left\{\mathbf{s}_{\rm N}(t)\mathbf{s}_{\rm N}^{\mathrm{H}}(t)\right\}=\text{diag}\{g_{K_{1}+1},g_{K_{1}+2},\cdots,g_{K}\}$ represent the source covariance matrices for the far- and near-field UAVs. In practice, the theoretical $\mathbf{R}$ is estimated by the sample average covariance matrix $\hat{\mathbf{R}}$, given by
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {+v_{i}(r_{k},\alpha_{k},\beta_{k})-v_{j}(r_{k},\alpha_{k},\beta_{k})\Big{)}}\!\!+\!\!\sigma^{2}\delta_{i,j},$
where $\delta_{i,j}=1$ when $i=j$, and $\delta_{i,j}=0$ otherwise. It is observed that the anti-diagonal entries of $\mathbf{R}$ only involve the angle information of both far- and near-field UAVs, which is given by
=${}_{n,N+1-n}\!=\!\!\sum_{k=1}^{K}\!g_{k}e^{j\frac{2\pi}{\lambda}2d_{0}(n_{x}\alpha_{k}+n_{y}\beta_{k})}\!\!+\!\!\sigma^{2}\delta_{n,N\!+\!1\!-\!n},n\in\mathcal{N},$ (21)
which endows the possibility to estimate the angles of UAVs.
@@
$[{\tilde{\bm{\eta}}}]_{n}=\sum_{k=1}^{K}g_{k}e^{j\frac{2\pi}{\lambda}2d_{0}(n_{x}\alpha_{k}+n_{y}\beta_{k})},\forall n\in\mathcal{N}.$ (22)
Then, we shall show that the entries of ${\tilde{\bm{\eta}}}$ can be used to construct a new matrix, which is equivalent to the covariance matrix of a virtual sparse UPA (S-UPA), hence allowing for applying subspace-based angle estimation methods.
Specifically, we consider a S-UPA with $\widetilde{N}\!\!=\!\widetilde{N}_{x}\widetilde{N}_{y}$ antennas where $\widetilde{N}_{x}\!=\!\frac{N_{x}+1}{2}$ and $\widetilde{N}_{y}\!=\!\frac{N_{y}+1}{2}$, and the inter-antenna spacing is $\lambda$. Then, the received signas at the virtual S-UPA is given by
@@
$[\widetilde{\mathbf{R}}_{i,j}]_{i_{x},j_{x}}=\sum\limits_{k=1}^{K}e^{\jmath\frac{4\pi}{\lambda}(i-j)\beta_{k}}e^{\jmath\frac{4\pi}{\lambda}(i_{x}-j_{x})\alpha_{k}},$ (24)
@@
Figure 3: Illustration of angle estimation. System parameters are set as $f=10$ GHz and $N_{x}=N_{y}=61$. The elevation and azimuth angles for far-field UAVs are ($\frac{\pi}{3}$, $-\frac{\pi}{4}$, $1000$ m) and ($\frac{\pi}{8}$, $\frac{2\pi}{3}$, $1500$ m), while the locations for near-field UAVs are set as ($\frac{5\pi}{13}$, ${0.23\pi}$, $50$ m) and ($\frac{\pi}{4}$, $\frac{5\pi}{21}$, $60$ m).
@@
The entries of ${\tilde{\bm{\eta}}}$ in (22) can be used to construct the matrix $\widetilde{\mathbf{R}}$. Mathematically, each entry of $[\widetilde{\mathbf{R}}]_{i,j}$ is given by
@@
.$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \big{.}[\tilde{\mathbf{y}}]_{\frac{N+1}{2}\!+\!(j-1)N_{x}\!:\!\frac{N+N_{x}}{2}\!+\!(j-1)N_{x}}\big{)},i,j\in\widetilde{\mathcal{N}}_{y}.$
Given Lemma 3, the 2D-angular MUSIC algorithm can be utilized to estimate UAVs’ angles by leveraging eigenvectors of the covariance matrix $\widetilde{\mathbf{R}}_{{\tilde{\bm{\eta}}}}$, given by
@@
where $\widetilde{\mathbf{U}}_{s}\in\mathbb{C}^{\widetilde{N}\times K}$ and $\widetilde{\mathbf{U}}_{n}\in\mathbb{C}^{\widetilde{N}\times\widetilde{N}-K}$ denote the signal and noise subspaces, respectively. In addition, $\widetilde{\bm{\Sigma}}_{s}$ and $\widetilde{\bm{\Sigma}}_{n}$ are diagonal matrices, representing the corresponding eigenvalues of the signal and noise subspaces, respectively. To be specific, based on the 2D MUSIC principle, the azimuth and elevation angles of both far-field and near-field UAVs can be estimated by searching peaks in the 2D-angular MUSIC spectrum [17]
@@
However, it is worth noting that this method introduces two issues in UAV angle estimation. First, there exist angular ambiguity in the angle estimation in (27), due to the sparsity of the virtual S-UPA, i.e., $\tilde{\mathbf{a}}^{H}_{\rm SUPA}(\alpha,\beta)=\tilde{\mathbf{a}}^{H}_{\rm SUPA}(\alpha\pm 1,\beta\pm 1)$. Second, performing a grid search in the 2D-angular domain incurs prohibitively high computational complexity. These issues will be addressed in the sequel.
@@
[…]
@@
$\begin{array}[]{ll}(\textbf{\text{P2}})&\min\limits_{\alpha,\beta}\leavevmode\nobreak\ \leavevmode\nobreak\ \tilde{\mathbf{a}}_{x}^{H}(\alpha)\widetilde{\mathbf{T}}(\beta)\tilde{\mathbf{a}}_{x}(\alpha)\\ &\leavevmode\nobreak\ \text{s.t.}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \tilde{\mathbf{c}}_{{\widetilde{N}_{x}}}^{{H}}\tilde{\mathbf{a}}_{x}(\alpha)=1,\end{array}$ (29)
[…]
@@
[…]

### v2

@@ II-B Signal Model
The $K$ targets are assumed to simultaneously transmit pilots in $M$ time slots which are in narrow-band and uncorrelated. As such, the received signal in the UPA antennas at time $m$, denoted by $\bm{\eta}(m)\in\mathbb{C}^{N\times 1}$, is given by
@@
where $\mathbf{G}\triangleq[\mathbf{A},\mathbf{B}]\in\mathrm{C}^{N\times K}$ denotes the array steering matrix (ASM) with $\mathbf{A}\triangleq[\mathbf{a}\left(\theta_{1},\phi_{1}\right),\mathbf{a}\left(\theta_{2},\phi_{2}\right),\cdots,\mathbf{a}\left(\theta_{K_{1}},\phi_{K_{1}}\right)]$ and $\mathbf{B}\triangleq[\mathbf{b}\left(r_{K_{1}+1},\theta_{K_{1}+1},\phi_{K_{1}+1}\right),\cdots,\mathbf{b}\left(r_{K},\theta_{K},\phi_{K}\right)]\in\mathbb{C}^{N\times K_{2}}$ denoting the ASM for the far- and near-field targets, respectively. The noise vector is distributed as $\mathbf{n}(m)\sim\mathcal{C\mathcal{N}}\left(0,\sigma^{2}\mathbf{I}_{M}\right)$ with zero mean and variance $\sigma^{2}$. Moreover, $\mathbf{s}(m)=[\mathbf{s}^{T}_{\rm F}(m),\mathbf{s}^{T}_{\rm N}(m)]^{T}\in\mathbb{C}^{K\times 1}$ denotes the source vector (i.e., transmitted signals by targets) with $\mathbf{s}_{\rm F}(m)=[s_{1}(m),s_{2}(m),\cdots,s_{K_{1}}(m)]^{T}\in\mathbb{C}^{K_{1}\times 1}$ and $\mathbf{s}_{\rm N}(m)=[s_{K_{1}+1}(m),s_{K_{1}+2}(m),\cdots,s_{K}(m)]^{T}\in\mathbb{C}^{K_{2}\times 1}$ representing the source sub-vectors for far- and near-filed targets, satisfying
$\mathbb{E}\{s_{j}(m)s_{k}(m)\}=\left\{\begin{matrix}g_{k},\penalty 10000\ \penalty 10000\ j=k,\\ 0,\penalty 10000\ \penalty 10000\ j\neq k,\end{matrix}\right$. (9)
where $p_{k}$ and $g_{k}=N|\beta_{k}|^{2}p_{k}$ denote the transmit power and equivalent channel power gain for the $k$-th target, respectively.
@@
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ =\mathbf{W}(m)\mathbf{A}\mathbf{s}_{{\rm F}}(m)\!\!+\!\!\mathbf{W}(m)\mathbf{B}\mathbf{s}_{{\rm N}}(m)\!\!+\!\!\mathbf{W}(m)\mathbf{n}(m),$
@@
where $\bm{\Pi}\in\mathbb{C}^{N\times N}$ and $\mathbf{v}_{f}(m)\in\mathbb{C}^{U\times 1},f\in\mathcal{F}\triangleq\{1,2,\cdots,N_{\rm RF}\}$ denote the permutation matrix determined by the sub-connected configuration and the combining vector of the $f$-th RF chain, respectively. For example, consider a UPA with $N_{x}=2$ and $N_{y}=4$ with $N_{\rm RF}=2$ RF chains. Following the antennas’ indices, the combining vectors of the first and second RF chains are given by $\mathbf{W}^{T}(m)]_{:,1}=[v_{1,1}(m),v_{1,2}(m),0,0,v_{1,3}(m),v_{1,4}(m),0,0]^{T}$, and $\mathbf{W}^{T}(m)]_{:,2}=[0,0,v_{2,1}(m),v_{2,2}(m),0,0,v_{2,3}(m),v_{2,4}(m)]^{T}$, respectively. In addition, the permutation matrix $\bm{\Pi}$ can be constructed as $\bm{\Pi}=[\mathbf{e}_{1},\mathbf{e}_{2},\mathbf{e}_{5},\mathbf{e}_{6},\mathbf{e}_{3},\mathbf{e}_{4},\mathbf{e}_{7},\mathbf{e}_{8}]^{T}$, where $\mathbf{e}_{n}$ denotes the $n$-th column of the identity matrix $\mathbf{I}_{8}$. Hence, the combining matrix $\mathbf{W}^{T}(m)$ can be rewritten as $\mathbf{W}^{T}(m)=\bm{\Pi}\times\text{diag}\{\mathbf{v}_{1}(m),\mathbf{v}_{2}(m)\}$, where $\mathbf{v}_{1}(m)=[v_{1,1}(m),v_{1,2}(m),v_{1,3}(m),v_{1,4}(m)]^{T}$ and $\mathbf{v}_{2}(m)=[v_{2,1}(m),v_{2,2}(m),v_{2,3}(m),v_{2,4}(m)]^{T}$.
@@ III Proposed Mixed-field Localization Method
In this section, we first elaborate the main design challenges when the BS of a hybrid antenna array is used for localizing the mixed-field targets. Then an efficient localization algorithm is proposed to solve them.
Specifically, compared with existing works on mixed-field localization, the studied problem differs in the following three main aspects, which introduce new challenges. First, most existing works on near-field localization method using decoupled subspace methods (e.g., [24]) rely on the assumption that the inter-antenna spacing is no larger than $\frac{\lambda}{4}$, such that reduced-dimensional/modified MUSIC algorithms can be devised to enable near-field localization. However, this assumption is incompatible to legacy wireless communication systems with typically half-wavelength inter-antenna spacing. Second, existing angle/location estimation methods for near-field localization primarily focus on digital array systems, which, however, may not be applicable to hybrid antenna systems with reduced-dimensional received signals. [^1] Although the beamspace MUSIC method was designed to effectively estimate path directions hybrid antenna systems [5], it inevitably decreases the maximum number of estimable sources $\mathcal{O}(N_{\rm RF})$. Third, many mixed-field localization methods typically differentiate far-field and near-field targets by detecting far-field targets’ angles in the angle-domain MUSIC spectrum by letting $r\to\infty$. However, it will be shown later in Section III-C that this method may make false detection of near-field targets, when their ranges are relatively large (still smaller than Rayleigh distance), hence resulting in high localization errors. This issue becomes more severe when the half-wavelength inter-antenna spacing constraint is considered due to the appearance of ambiguous targets’ angles.
@@
  • In Step 1, we accurately recover equivalent received signals at the UPA antennas from the reduced-dimensional signals $\mathbf{y}(m)$, by judiciously designing the analog combining matrices over time $\{\mathbf{W}(m)\}$. In particular, the UPA with half-wavelength inter-antenna spacing is considered, which inevitably causes angular ambiguity and will be addressed in the third step.
  • In Step 2, we design a modified low-dimensional MUSIC algorithm for UPA-based target localization by constructing a new Toeplitz-structured (far-field equivalent) covariance matrix of a virtual S-UPA, which enables effective decoupling of the angle and range estimation.
  • In Step 3, we propose an efficient target classification method based on distinct characteristics of 3D MUSIC spectrum in the range domain between far-field and near-field targets, which can also effectively address the angular ambiguity issue caused by the virtual S-UPA arsing from the half-wavelength inter-antenna spacing the physical UPA.
@@ III-A Signal Reconstruction
To this end, we divide the $M$ received signals $\mathbf{y}(m)$ into $L=M/U$ groups, for which the $\ell$-th group’s signals can be concatenated into a new effective signal vector $\mathbf{Y}_{\ell}\in\mathbb{C}^{N\times 1}$ as given in (12). As such, each group’s received pilots can be utilized to recover a sample of equivalent signals received at the UPA. [^2] In (12), for ease of design, the transmitted signals from targets are assumed to be identical in every $M$ pilots, i.e., $\mathbf{s}\big(M(\ell-1)+1\big)=\mathbf{s}\big(M(\ell-1)+2\big)=,\cdots,=\mathbf{s}\big(M(\ell-1)+M\big)\triangleq\mathbf{s}_{\ell}$. Moreover, for simplicity, we set $\mathbf{W}(u)=\mathbf{W}\big(u+\ell U\big),\forall u\in\mathcal{U}\triangleq\{1,2,\cdots,U\},\forall\ell\in\mathcal{L}$. As such, if ${\mathbf{W}_{0}}\triangleq[\mathbf{W}^{T}(1),\mathbf{W}^{T}(2),\cdots,\mathbf{W}^{T}(U)]^{T}$ is an invertible matrix, the $\ell$-th sample of the reconstructed signals received at the UPA antennas can be recovered as [17]
@@
$({\bf P1}):\penalty 10000\ \penalty 10000\$$\min_{\mathbf{W}_{0}}\penalty 10000\ {\rm Tr}\Big(\mathbf{n}_{\ell}\mathbf{n}_{\ell}^{H}\Big)$ (15a)
$\penalty 10000\ {\text{s}}{\text{.t}}{\rm{.}}\penalty 10000\ \penalty 10000\ \penalty 10000\ {\rm Rank}(\mathbf{W}_{0})=N,$ (15b)
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ |[\mathbf{v}_{f}(u)]_{i}|=1,\forall u,i\in\mathcal{U},\forall f\in\mathcal{F},$ (15c)
@@
Although Problem (P1) is non-convex, the necessary conditions for its optimal solution is given as follows.
@@ III-B Angle Estimation
In this subsection, we develop a decoupled MUSIC algorithm tailored to the UPA architecture. Owing to the half-wavelength spacing constraint, it is difficult to directly construct an ambiguity-free equivalent far-field covariance matrix. To address this issue, we first construct a far-field covariance matrix of a virtual sparse UPA (S-UPA) from the original covariance matrix, which enables the decoupling of the joint 3D angle-and-range estimation into separate angle and range estimations. Based on the resulting angle-only covariance matrix of the virtual S-UPA, a decoupled 2D angular MUSIC algorithm is further devised to separately estimate the elevation and azimuth angles. The sparse structure of the virtual S-UPA inevitably introduces angular ambiguities, which will be effectively resolved in Section III-C.
@@
where $\mathbf{S}_{\rm F}=\mathbb{E}\left\{\mathbf{s}_{\rm F}(t)\mathbf{s}_{\rm F}^{\mathrm{H}}(t)\right\}=\text{diag}\{g_{1},g_{2},\cdots,g_{K_{1}}\}$ and $\mathbf{S}_{\rm N}=\mathbb{E}\left\{\mathbf{s}_{\rm N}(t)\mathbf{s}_{\rm N}^{\mathrm{H}}(t)\right\}=\text{diag}\{g_{K_{1}+1},g_{K_{1}+2},\cdots,g_{K}\}$ represent the source covariance matrices for the far- and near-field targets. In practice, the theoretical $\mathbf{R}$ is estimated by the sample average covariance matrix $\hat{\mathbf{R}}$, given by
@@
$\penalty 10000\ \penalty 10000\ \penalty 10000\ {+v_{i}(r_{k},\alpha_{k},\beta_{k})-v_{j}(r_{k},\alpha_{k},\beta_{k})\Big)}\!\!+\!\!\sigma^{2}\delta_{i,j},$
where $\delta_{i,j}=1$ when $i=j$, and $\delta_{i,j}=0$ otherwise. It is observed that the anti-diagonal entries of $\mathbf{R}$ only involve the angle information of both far- and near-field targets, which is given by
${}_{n,N+1-n}\!=\!\!\sum_{k=1}^{K}\!g_{k}e^{j\frac{2\pi}{\lambda}2d_{0}(n_{x}\alpha_{k}+n_{y}\beta_{k})}\!\!+\!\!\sigma^{2}\delta_{n,N\!+\!1\!-\!n},n\in\mathcal{N},$ (21)
which endows the possibility to estimate the angles of targets.
@@
${\tilde{\bm{\eta}}}]_{n}=\sum_{k=1}^{K}g_{k}e^{j\frac{2\pi}{\lambda}2d_{0}(n_{x}\alpha_{k}+n_{y}\beta_{k})},\forall n\in\mathcal{N}.$ (22)
Then, we shall show that the entries of $\tilde{\bm{\eta}}$} can be used to construct a new matrix, which is equivalent to the covariance matrix of a virtual S-UPA, hence allowing for applying subspace-based angle estimation methods.
Specifically, we consider an S-UPA with $\widetilde{N}\!\!=\!\widetilde{N}_{x}\widetilde{N}_{y}$ antennas where $\widetilde{N}_{x}\!=\!\frac{N_{x}+1}{2}$ and $\widetilde{N}_{y}\!=\!\frac{N_{y}+1}{2}$, and the inter-antenna spacing is $\lambda$. Then, the received signas at the virtual S-UPA is given by
@@
$\widetilde{\mathbf{R}}_{i,j}]_{i_{x},j_{x}}=\sum\limits_{k=1}^{K}e^{\jmath\frac{4\pi}{\lambda}(i-j)\beta_{k}}e^{\jmath\frac{4\pi}{\lambda}(i_{x}-j_{x})\alpha_{k}},$ (24)
@@
Figure 3: Illustration of angle estimation. System parameters are set as $f=10$ GHz and $N_{x}=N_{y}=61$. The elevation and azimuth angles for far-field targets are ($\frac{\pi}{3}$, $\frac{\pi}{4}$, $1000$ m) and ($\frac{\pi}{8}$, $\frac{2\pi}{3}$, $1500$ m), while the locations for near-field targets are set as ($\frac{5\pi}{13}$, $0.23\pi$}, $50$ m) and ($\frac{\pi}{4}$, $\frac{5\pi}{21}$, $60$ m).
@@
The entries of $\tilde{\bm{\eta}}$} in (22) can be used to construct the matrix $\widetilde{\mathbf{R}}$. Mathematically, each entry of $\widetilde{\mathbf{R}}]_{i,j}$ is given by
@@
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \big.[\tilde{\mathbf{y}}]_{\frac{N+1}{2}\!+\!(j-1)N_{x}\!:\!\frac{N+N_{x}}{2}\!+\!(j-1)N_{x}}\big),i,j\in\widetilde{\mathcal{N}}_{y}.$
Given Lemma 3, the 2D-angular MUSIC algorithm can be utilized to estimate targets’ angles by leveraging eigenvectors of the covariance matrix $\widetilde{\mathbf{R}}_{{\tilde{\bm{\eta}}}}$, given by
@@
where $\widetilde{\mathbf{U}}_{s}\in\mathbb{C}^{\widetilde{N}\times K}$ and $\widetilde{\mathbf{U}}_{n}\in\mathbb{C}^{\widetilde{N}\times\widetilde{N}-K}$ denote the signal and noise subspaces, respectively. In addition, $\widetilde{\bm{\Sigma}}_{s}$ and $\widetilde{\bm{\Sigma}}_{n}$ are diagonal matrices, representing the corresponding eigenvalues of the signal and noise subspaces, respectively. To be specific, based on the 2D MUSIC principle, the azimuth and elevation angles of both far-field and near-field targets can be estimated by searching peaks in the 2D-angular MUSIC spectrum [6]
@@
However, it is worth noting that this method introduces two issues in target angle estimation. First, there exist angular ambiguity in the angle estimation in (27), due to the sparsity of the virtual S-UPA, i.e., $\tilde{\mathbf{a}}^{H}_{\rm SUPA}(\alpha,\beta)=\tilde{\mathbf{a}}^{H}_{\rm SUPA}(\alpha\pm 1,\beta\pm 1)$. Second, performing a grid search in the 2D-angular domain incurs prohibitively high computational complexity. These issues will be addressed in the sequel.
@@
To this end, we first rewrite the denominator of $F(\alpha,\beta)$ in (27) as
@@
$\begin{array}[]{ll}(\textbf{\text{P2}})&\min\limits_{\alpha,\beta}\penalty 10000\ \penalty 10000\ \tilde{\mathbf{a}}_{x}^{H}(\alpha)\widetilde{\mathbf{T}}(\beta)\tilde{\mathbf{a}}_{x}(\alpha)\\ &\penalty 10000\ \text{s.t.}\penalty 10000\ \penalty 10000\ \penalty 10000\ \tilde{\mathbf{c}}_{{\widetilde{N}_{x}}}^{{H}}\tilde{\mathbf{a}}_{x}(\alpha)=1,\end{array}$ (29)
where $\tilde{\mathbf{c}}_{{\widetilde{N}_{x}}}^{T}=[\underbrace{0,\cdots,0,}_{\frac{\widetilde{N}_{x}-1}{2}}1,\underbrace{0,\cdots,0}_{\frac{\widetilde{N}_{x}-1}{2}}]$ and the constraint of $(\textbf{\text{P2}})$ is to avoid the trivial solution $\tilde{\mathbf{a}}_{x}(\alpha)=\mathbf{0}$. Although it is not easy to obtain the optimal solution to Problem (P2) in closed form, we present its poperies in the lemma below.
@@
Proof: We first construct a Lagrange function $L(\alpha,\beta)$ for $(\textbf{\text{P2}})$, which is given by